Recent Maintenance – KDAK

 KDAK operational status:

2017.11.25

Firmware for both Afar radios was upgraded from v7.19 to v7.21 on 11/25 (2017,329-09:20).

2017.10.27

An IDA engineer visited the station and replaced the Episensor strong motion sensor and installed a paroscientific microbarometer.

2017.08.21

IDA engineers visited the station and replaced the KS54000 seismometer with a Trillium 360PH seismometer, performed maintenance on the TEG, and radios.

2017.01.27

The vertical channels from the KS54000 seimometer are seismic as of day 2016,352(Dec 17) and the horizontal channels are seismic as of day 2016,358.

2016.12.06

The station is back up as of day 2016:341-22:02 after the FPC board for the TEG was replaced.  All channels from the KS54000 seismometer are non-seismic since power was restored for reasons yet to be determined.

2016.11.21

The station is down as of day 2016:320-23:42:44 due to problems with the TEG.

2016.09.23

IDA engineers visited the station and installed a Vaisala weather transmitter, Hyperion infrasound sensor, new station computers, and performed maintenance on the TEG.

2016.08.05

The GPS clocks at KDAK were unlocked on day 2016,210-01:56 due to damage to the GPS antennas caused by bears.  The station was also down from day 2016,216-12:19 to 2016,217-21:15 due to running out of propane for the TEG tanks.  Data acquisition was restored with locked GPS clocks after the station operator replaced the GPS antennas and the propane tanks were filled.

2015.11.24

The station is back online as of 2015,328-23:40 after the station operator powered up the TEG and restored power.  The station power was down due to running out of propane for the TEG tanks.

2015.11.23

The station is down as of day 2015,324-02:13 due to power problems. IDA engineers are working with the station operator to resolve this issue.

2015.10.08

Missing

2015.10.05

The station is down as of 2015,277-22:00 due to communication problems to the wellhead.  The station operator has been requested to investigate.

2015.09.10

IDA engineers relocated the GSN station computers and the CTBTO interface to the wellhead to accommodate PTS requirements.  To compensate for the increased power load, they upgraded the thermoelectric generator (TEG) at the same time.

2014.10.08

An IDA engineer has replaced the STS-2 with a new Trillium 120PH seismometer, and replaced the Setra barometer.

2014.02.25

The station was down from day 2014,055-07:42 to 2014,056-15:59 due to a power outage. Data acquisition was restored after the station operator visited the station and cycled power to the UPS and rebooted the station computers.

2013.07.11

Telemetry was restored after the station operator cycled power to the cell phone modem.

2013.07.09

Telemetry to the station has been down since day 2013,187-17:42 due to problems with the modem that are yet to be determined. The station operator has been instructed to cycle power to the modem.

2012.12.11

An IDA engineer is on-site and has restored telemetry to the station. Equipment installed include a cell phone modem, episensor strong motion sensor, and a barometer.

2012.10.23

The station operator reports that the local transformer has been replaced and that normal power to the station has been restored. IDA engineers are currently working with the station operator and internet service provider to restore the telemetry connection to the station.

2012.09.25

Telemetry to IDA station KDAK is down due to a power outage. The station is still expected to be acquiring data since wellhead equipment is running off of a generator.

2011.09.15

The North-South channels from the FBA-23 strong motion sensor are spiking excessively since day 2011,258-20:29.

2011.07.13

The output of the KS-54000 North-South channel has been variable and anomalously low since day 2009,223 for unknown causes believed to be related to the sensor’s downhole electronics.

2011.07.04

An IDA engineer visited the station and upgraded it to a Quanterra Q330HR system. Seismometers include KS-54000 and STS-2 seismometers, and an FBA-23 strong motion sensor.

2010.09.22

The GPS clock at KDAK is locked as of day 2010,264-23:29 after the station operator cylced power to the GPS clock.

2010.09.19

The GPS clock at KDAK is in an unlocked state as of day 2010,262-06:43. IDA engineers are corresponding with the station operator to resolve this problem.

2009.09.17

The station operator replaced the microbarometer and restored the WDO channel.

2009.08.13

An IDA engineer visited the station to move equipment, as there is a new water treatment system being installed at the site. While on-site there were problems with the GPS clock, the DAS, and telemetry to the station. The engineer replaced the GPS clock with one shipped from UCSD, and replaced a faulty circuit board in the DAS. Another IDA engineer worked with the local leased line provider to restore telemetry. There are currently problems with the microbarometer. IDA engineers are troubleshooting the problem with the help of the station operator.

2008.08.13

The WDO channel has been restored as of about day 2008,225-03:07 after the station operator replaced the microbarometer.

2008.02.02

The station was down from about day 2008,034 to 2008,036 due to a start-up problem in the DAS after the station operator performed maintenance on the power system. The station operator restored data acquisition by cycling power to the DAS.

2007.01.29

The station was operational from day 026-21:12 to 029-18:20. Power supply problems persisted as the station was going off and on line intermittently. Normal data acquisition has resumed since about day 030.

2007.01.26

The WDO-00 data from the microbarometer are flatlined as of day 026-21:11 for reasons yet to be determined.

2007.01.18

IDA station KDAK is down as of about day 018-12:40 due to a power supply problem. The station operator has been requested to look for a spare converter limiter module for the generator.

2005.04.11

The station is up as of 04/09 after the generator powering the DAS was fixed.

2005.03.28

The station is down due as of 03/22 due to problems with power to the station.

2004.08.09

Station operating normally.

2003.09.08

Temporary data gap from 8/30 to 9/5 due to workstation shutdown. Data acquired again after power to workstation restored.

2003.06.13

Power failure. Power was out longer than UPS battery back-up and station equipment temporarily shut down.

2003.04.09

New routers at station and gateway, linked via leased line connection.

2002.10.16

Upgrade of data acquisition system to MK8 is complete.

2002.10.08

Data acquisition is currently down while the station’s data acquisition system is upgraded.

2002.04.22

All channels from the FBA-23 are dead. Bursts of low frequency noise on KS-54000 vertical channels noted.